A car travelling at 40 km/hr accelerates at 2.3 m/s2 for 2.7 s.  How far has it travelled in that time?   What is its final velocity?


1
Expert's answer
2021-02-11T10:40:55-0500

a)

"d=v_0t+\dfrac{1}{2}at^2,""d=11.1\ \dfrac{m}{s}\cdot2.7\ s+\dfrac{1}{2}\cdot2.3\ \dfrac{m}{s^2}\cdot(2.7\ s)^2=38.35\ m."

b)

"v=v_0+at,""v=11.1\ \dfrac{m}{s}+2.3\ \dfrac{m}{s^2}\cdot2.7\ s=17.31\ \dfrac{m}{s}."